Suppose a consumer having a disposable income of 300 birr consumers only two commodities X and Y and his utility function is given as U(X,Y)=50X-X2+25Y given further that price of X and Y are 1 birr and 2.5 birrr respectively, determine:

A. The consumer's optimal bundles

B. His Marginal utility of income and its interpretation


1
Expert's answer
2022-02-14T15:19:30-0500

"U_(x,y)=50X-2X+25Y"

Budget line:

"300=X+2.5Y"

At optimal bundle, slope of the budget line is equal to the slope of Indifference curve.

Slope of BL"=\frac{-y}{x}=\frac{-2.5}{1}=-2.5"

Slope of IC"=\frac{-MU_x}{MU_y}"

"MU_x=48+25Y"

"MU_y=48X+25"

Therefore slope of IC"=\frac{-(48+25Y)}{48X+25}"

Equating slope of BL and IC:

"\frac{-(48+25Y)}{48X+25}=\frac{-2.5}{1}"

"-48-25Y=120X-62.5"

"25Y=120X+14.5"

"Y=4.8X+0.58"

Replacing Y in the budget line:

"300=X+2.5\u00d7(4.8X+0.58)"

"300=13X+1.45"

"13X=298.55"

"X=22.97"

"Y=4.8X+0.58"

"Y=4.8\u00d722.97+0.58=110.84"

B)

"U_{x,y}=48X+25Y"

"=48\u00d722.97-2\u00d722.97+25\u00d7110.84"

"=3873.56"